Solve the system of equations. (If the system is dependent, enter a general solution in terms of c. If there is no solution, enter NO SOLUTION.)
x + 3y - 4z = 17
4x - y + z = 17
3x + 4y - 3z = 24
(x, y, z)=

Answer by TimothyLamb(4379)

x + 3y - 4z = 17
4x - y + z = 17
3x + 4y - 3z = 24
---
put the system of linear equations into standard form
---
x + 3y - 4z = 17
4x - y + z = 17
3x + 4y - 3z = 24
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 5
y = 0
z = -3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php